# Integrated Push-and-Pull Update Model for Goal-Oriented Effective Communication

### v1

@@ IV-D1 Computing $\pi_{\gamma,\mu}^{*}$
$V_{k}^{\pi_{\gamma,\mu}}(s)$ $=\mathbb{E}\big{[}r_{k}+\lambda r_{k+1}+\lambda^{2}r_{k+2}+\cdot\cdot\cdot~{}|~{}s_{k}=s\big{]}$
$\approx\mathbb{E}\big{[}r_{k}+\lambda V_{k-1}^{\pi_{\gamma,\mu}}~{}|~{}s_{k}=s\big{]}$ (17)
@@
which takes into account the cost caused by the taken action.
@@
by using the span seminorm [23, Section 6.6.1]. As the decision policies are unichain and have aperiodic transition matrices, the criterion in (21) is satisfied after finite iterations for any value of $\lambda\in[0,1]$ [23, Theorem 8.5.4].
@@ IV-D2 Computing $\mu^{*}$
We leverage the bisection search method to compute the optimal Lagrange multiplier over multiple steps in the outer loop based on the derived $\pi_{\gamma,\mu}^{*}$ from the inner loop. Starting with an initial interval $[\mu^{-},\mu^{+}]$ such that $h_{\gamma}(\mu^{-})h_{\gamma}(\mu^{+})<0$, the value of the multiplier at the $l$-th, $\forall l\in\mathbb{N}$, step is improved by $\mu^{(l)}=\frac{\mu^{-}+\mu^{+}}{2}$. As shown in Algorithm 1, at each step, the value of either $\mu^{-}$ or $\mu^{+}$ and the corresponding decision policy $\pi_{\gamma,\mu^{-}}$ or $\pi_{\gamma,\mu^{+}}$, respectively, are updated according to the cost constraint in (IV-B3) until a stopping criterion $\rvert\mu^{+}-\mu^{-}\rvert<\varepsilon_{\mu}$ is reached with the accuracy $\varepsilon_{\mu}$. Considering (IV-C) and (14), $h_{\gamma}(\mu)$ is a linear non-increasing function of $\mu$. In this regard, the bisection method searches for the smallest Lagrange multiplier that guarantees the cost constraint. Also, one can show that $h_{\gamma}(\mu)$ denotes a Lipschitz continuous function with the Lipschitz constant as below
$\bigg{\rvert}C_{\gamma,\rm max}-\underset{N\rightarrow\infty}{\lim\sup}~{}\frac{1}{N}\,\mathbb{E}\bigg{[}\sum_{n=1}^{N}\lambda^{n}c_{\gamma}(\gamma_{n})\bigg{]}\bigg{\rvert}.$
Therefore, the bisection search converges to the optimal value of $\mu$ within $L\in\mathbb{N}$ finite steps [25, pp. 294].
@@
$\mathcal{C}:~{}\underset{N\rightarrow\infty}{\lim\sup}~{}\frac{1}{N}\,\mathbb{E}\bigg{[}\sum_{n=1}^{N}\lambda^{n}c_{\gamma}(\gamma_{n})\bigg{]}=C_{\gamma,\rm max}.$ (23)
Otherwise, the derived policy becomes randomized stationary in the shape of mixing two deterministic policies $\pi_{\gamma,\mu^{-}}=\underset{\mu\rightarrow\mu^{-}}{\lim}\pi_{\gamma,\mu}$ and $\pi_{\gamma,\mu^{+}}=\underset{\mu\rightarrow\mu^{+}}{\lim}\pi_{\gamma,\mu}$ with probability $\eta\in[0,1]$ [26]. Hence, we can write
$\pi_{\gamma}^{*}\leftarrow\eta\pi_{\gamma,\mu^{-}}+(1-\eta)\pi_{\gamma,\mu^{+}},$ (24)
which implies that the decision policy is randomly chosen as $\pi^{*}_{\gamma}=\pi_{\gamma,\mu^{-}}$ and $\pi^{*}_{\gamma}=\pi_{\gamma,\mu^{+}}$ with probabilities $\eta$ and $1-\eta$, respectively. In (24), $\eta$ is computed such that the condition $\mathcal{C}$ in (23) is maintained.
@@ IV-D3 Complexity analysis
The value iteration approach in the inner loop is polynomial with $\mathcal{O}(\rvert\mathcal{A}_{\gamma}\rvert\rvert\mathcal{S}_{\gamma}\rvert^{2})$ arithmetic operations at each iteration. Thereby, the longest running time of Algorithm 1, in terms of the number of arithmetic operations over both loops, is given by $\mathcal{O}\!\left(\frac{2L\rvert\mathcal{S}_{\gamma}\rvert^{2}}{1-\lambda}\log\!\left(\frac{1}{1-\lambda}\right)\!\right)$ for the fixed $\lambda$, as studied in [27] and [28]. The complexity of the algorithm increases with a larger state space, additional steps in the outer loop, and as $\lambda\rightarrow 1$.
@@ V Monte Carlo Probability Distribution Estimation
Within the E-horizon, the SA does not make any decisions. Instead, it focuses on communicating updates at the highest possible rate while adhering to cost constraints. Once receiving these updates, the AA measures their usefulness, stores them in memory, and sends E-ACK signals for effective updates. The SA logs whether the E-ACK has been successfully received or not in every slot of the E-horizon. Finally, employing the received E-ACK signals at the SA and the measured updates’ usefulness at the AA, both agents perform their estimations.
[…]
[scale=0.86]Media/Timing.eps
@@ V-A Usefulness Probability of Received Updates
Picking the $j$-th, $\forall j\in\mathcal{J}$, outcome from the set $\hat{\mathcal{V}}$ (defined in Section III-A) that corresponds to the received update’s usefulness from the endpoint’s perspective in the $m$-th slot of the E-horizon, i.e., $\hat{v}_{m}$, the relevant estimated probability of that outcome is given by
@@ V-B Probability of the Mapped Target Usefulness
We assume that the mapped target usefulness, i.e., $\hat{v}_{\rm tgt}$, is a member of the set $\hat{\mathcal{V}}_{\rm tgt}=\{\vartheta_{j}~{}|~{}j\in\mathcal{J}_{\rm tgt}\}$ with i.i.d. elements, where $\mathcal{J}_{\rm tgt}=\{1,2,...,\lvert\hat{\mathcal{V}}_{\rm tgt}\rvert\}$, and the probability of the $j$-th element is equal to $p_{\hat{v}_{\rm tgt}}(\hat{v}_{\rm tgt}=\vartheta_{j})$. Herein, as mentioned earlier, $p_{\hat{v}_{\rm tgt}}(\cdot)$ is the estimated pmf of the mapped target usefulness and obtained by
$p_{\hat{v}_{\rm tgt}}(\vartheta_{j})$ $=\sum_{e\in\{0,1\}}p_{\hat{v}_{\rm tgt}}\big{(}\vartheta_{j}~{}|~{}\hat{E}=e\big{)}\mathrm{Pr}\big{(}\hat{E}=e\big{)}$ (26)
@@
where $\hat{E}_{m}$ indicates the E-ACK arrival status in the $m$-th slot of the E-horizon. Furthermore, to derive the conditional probability in (26), we first consider the successful arrivals of E-ACK signals such that
$p_{\hat{v}_{\rm tgt}}\big{(}\vartheta_{j}~{}|~{}\hat{E}=1\big{)}$ $=\frac{\sum_{i\in\mathcal{I}}p_{\nu|\hat{E}=1}\big{(}\nu_{i}~{}|~{}\nu_{i}\geq\vartheta_{j}\big{)}}{\sum_{j\in\mathcal{J}_{\rm tgt}}\sum_{i\in\mathcal{I}}p_{\nu|\hat{E}=1}\big{(}\nu_{i}~{}|~{}\nu_{i}\geq\vartheta_{j}\big{)}}.$ (28)
@@
$p_{\hat{v}_{\rm tgt}}\big{(}\vartheta_{j}~{}|~{}\hat{E}=0\big{)}$ $=\frac{\sum_{i\in\mathcal{I}}p_{\nu|\hat{E}=0}\big{(}\nu_{i}~{}|~{}\nu_{i}<\vartheta_{j}\big{)}}{\sum_{j\in\mathcal{J}_{\rm tgt}}\sum_{i\in\mathcal{I}}p_{\nu|\hat{E}=0}\big{(}\nu_{i}~{}|~{}\nu_{i}<\vartheta_{j}\big{)}}.$ (29)
@@ VI-A Setup and Assumptions
TABLE I: Parameters for simulation results.
@@
Controlled transmission rate (value-agnostic) $0.8$
Controlled query rate (value-agnostic) $0.8$
@@ VI-B Results and Discussion
Fig. 5 illustrates the evolution of the average cumulative effectiveness over time for the push-and-pull model and different agent decision policies. As it is shown, applying the effect-aware policy in both agents offers the highest effectiveness, where the gap between its and the other policies’ performance increases gradually as time passes. Nevertheless, using the other value-agnostic policies at either or both agent(s) diminishes the effectiveness performance of the system by at least $12\%$ or $36\%$, respectively. Particularly, if the SA and the AA apply the effect-aware and Markovian effect-agnostic policies, sequentially, the offered effectiveness is even lower than the scenario in which both agents use periodic effect-agnostic policies. One of the reasons is that the Markovian query raising misleads the SA in estimating the mapped target usefulness of updates. It is also worth mentioning that throughout the E-horizon from slot $1$ to $1\times 10^{5}$, the scenarios where the AA raises effect-aware queries offer the same performance and better than those of the other scenarios. However, in the D-horizon, a performance gap appears and evolves depending on the applied policy at the SA. Therefore, by making effect-aware decisions at both agents based on estimations in the E-horizon and CMDP-based policies in the D-horizon, performance can be significantly improved.
[…]
[scale=0.4]Media/Effect_time.eps
@@
Figure 5: The evolution of the average effectiveness accumulated over time based on the push-and-pull model.
@@
The bar chart in Fig. 6 shows the average rate of update transmissions from the SA and the consequent actions performed at the AA based on raised queries that result in the system’s effectiveness, as depicted in Fig. 5. Interestingly, the scenario where both agents apply the effect-aware policy has lower transmission and action rates compared to the other scenarios except the ones with Markovian effect-agnostic query policies. However, the number of actions that can be done for those scenarios with Markovian queries is limited since most of the updates are received out of the action windows. Figs. 5 and 6 show that using effect-aware policies at both the SA and the AA not only brings the highest effectiveness but also needs lower update transmissions by an average of $11\%$, saving resources compared to the scenarios that have comparable performances. We also note that although effect-aware and periodic query decisions with effect-aware update transmission have almost the same action rate, the effect-aware case leads to more desirable effects or appropriate actions at the endpoint. This results in around $16\%$ higher average cumulative effectiveness, referring to Fig. 5.
[…]
[scale=0.4]Media/RateBar_v2.eps
@@
Figure 6: The average update transmission and query rates of different decision policies in the push-and-pull model.
[…]
[scale=0.4]Media/GradeEffect_time_PushPull.eps
[…]
[scale=0.4]Media/GradeEffect_time_Push.eps
[…]
[scale=0.4]Media/GradeEffect_time_Pull.eps
@@
Figure 7: The evolution of the average cumulative GoE over time, following the (a) push-and-pull, (b) push-based, and (c) pull-based models.
@@
Figs. 7(a), 7(b), and 7(c) present the average cumulative GoE provided in the system over $5\times 10^{5}$ time slots for the push-and-pull, push- and pull-based communication models, respectively. The corresponding effectiveness at the endpoint for the primary model can be found in Fig. 5. The plots demonstrate that when both agents decide based on effect-aware policies, regardless of the update communication model, the highest offered GoE of the system is reached. However, in other scenarios, the performance of some policies exceeds those of others, depending on the update model. For instance, having effect-aware update transmission and Markovian queries shows $2.52$ times higher average GoE for the push-based model than the push-and-pull one. This is because the push-based model has a larger action window.
@@
Besides, comparing Figs. 7(a) and 7(b), the reason that the provided GoE by the effect-aware decisions at both agents is $28\%$ lower for the push-based model compared to the push-and-pull is that the AA has to be available longer, which causes a higher cost. Also, with the pull-based model as in Fig. 7(c), the average GoE within a period is less than the average cost since the AA is only available to act at query instants, significantly reducing the average GoE despite the high update transmission rate. In the pull-based model, however, applying the CMDP-based update transmission decisions at the SA can address this issue with $16\%$ higher average GoE.
@@
The trade-off between the average effectiveness and the width of the action window, i.e., $\Theta_{\rm max}$, is shown in Fig. 8 for different decision policies. We see that the system cannot offer notable effectiveness with $\Theta_{\rm max}=1$, i.e., under the pull-based model. However, by expanding the action window width from $\Theta_{\rm max}=1$ to $10\,[\text{slot}]$, the average effectiveness boosts from its lowest to the highest possible value. Since all policies already reach their best performance before $\Theta_{\rm max}=10$, which indicates the push-based model, we can conclude that the push-and-pull model with a flexible action window is more advantageous than the push-based one with a very large fixed window. In addition, the scenario where both agents use effect-aware policies outperforms the others for $\Theta_{\rm max}>1$. It is worth mentioning that the performance of the scenarios where the AA raises Markovian queries converges to the highest level for very large widths. The average cumulative effectiveness of these scenarios, as in Fig. 5, visibly rises for $\Theta_{\rm max}\geq 9$.
[…]
[scale=0.4]Media/Effect_actionwindow.eps
@@
Figure 8: The average effectiveness versus the width of the action window for the push-and-pull model.
[…]
[scale=0.4]Media/Effect_sensitivity.eps
@@
Figure 9: The average effectiveness w.r.t. the target effectiveness grade in the push-and-pull model.
@@
Fig. 10 depicts the average effectiveness obtained in the system through $5\times 10^{5}$ time slots versus the controlled update transmission rate for the push-and-pull model and various agent decision policies. Concerning Section II-B, this controlled rate is related to the value-agnostic update transmission policies at the SA and denotes the expected number of updates to be communicated within the specified period of time. Therefore, the performance of the other policies should remain fixed for different controlled transmission rates. Fig. 10 reveals that increasing the controlled update rate increases the offered effectiveness when the SA applies the value-agnostic policies. However, even at the highest possible rate, subject to the maximum discounted cost, using effect-aware policies at both agents is necessary to ensure the highest average effectiveness, regardless of the controlled transmission rate. As an illustrative example, when the AA raises effect-aware queries, the effectiveness drops by an average of $38\%$ ($43\%$) if the SA transmits updates based on the periodic (Markovian) effect-agnostic policy instead of using the effect-aware one.
[…]
[scale=0.4]Media/Effect_transmitrate.eps
Figure 10: The performance comparison between various policies under different update transmission rates but fixed query rates for value-agnostic policies.
In Fig. 11, the plot shows the same trend as Fig. 10, but this time it focuses on the changes in the average effectiveness versus the controlled query rate. Herein, the controlled query rate is dedicated to the scenarios where the AA operates under effect-agnostic policies. The results indicate that the average effectiveness rises via the increase of the controlled query rate for the periodic and Markovian policies. However, the effectiveness increase is not significant for the latter one. Despite this, even with the highest rates, the effectiveness offered in the course of effect-aware queries is still higher than those of raising effect-agnostic queries. Therefore, the highest average effectiveness is achieved when both agents make effect-aware decisions, as depicted in Figs. 10 and 11.
[…]
[scale=0.4]Media/Effect_queryrate.eps
@@
Figure 11: The comparison between different policies with variable query rates but fixed update transmission rates for value-agnostic policies.
@@
In the context of the decision-making problem $\mathcal{P}_{2}$ in (IV-B3), altering the maximum discounted cost, i.e., $C_{\gamma,\rm max}$, can impact the decisions made by each agent. To study this, we have plotted Fig. 12 for the push-and-pull model under different decision policies. The figure shows that the stricter the cost constraint, the lower the average effectiveness, irrespective of the decision policy. Also, for all cost constraints, the scenario in which both agents use effect-aware policy yields the best performance, whereas the other policies outperform each other under different cost constraints. Due to CMDP-based decisions, the gap between the effectiveness performance of the best scenario and those of the others increases as the constraint decreases until $C_{\gamma,\rm max}=0.1$, where the SA can transmit all updates, and the AA can raise queries without restrictions.
[…]
[scale=0.4]Media/Effect_cost.eps
@@
Figure 12: The average effectiveness attained by different policies versus the maximum discounted cost in the push-and-pull model.

### v2

@@ IV-D1 Computing $\pi_{\gamma,\mu}^{*}$
$V_{k}^{\pi_{\gamma,\mu}}(s)$ $=\mathbb{E}\big{[}r_{k}+\lambda r_{k+1}+\lambda^{2}r_{k+2}+\cdot\cdot\cdot\,\lvert\,s_{k}=s\big{]}$
$\approx\mathbb{E}\big{[}r_{k}+\lambda V_{k-1}^{\pi_{\gamma,\mu}}\,\lvert\,s_{k}=s\big{]}$ (17)
@@
which takes into account the cost caused by the action taken.
@@
by using the span seminorm for the arbitrary $k^{\prime}$-th iteration [23, Section 6.6.1]. As the decision policies are unichain and have aperiodic transition matrices, the criterion in (21) is satisfied after finite iterations for any value of $\lambda\in[0,1]$ [23, Theorem 8.5.4].
@@ IV-D2 Computing $\mu^{*}$
We leverage the bisection search method to compute the optimal Lagrange multiplier over multiple steps in the outer loop based on the derived $\pi_{\gamma,\mu}^{*}$ from the inner loop. Starting with an initial interval $[\mu_{-}^{(0)},\mu_{+}^{(0)}]$ such that $h_{\gamma}(\mu_{-}^{(0)})h_{\gamma}(\mu_{+}^{(0)})<0$, the value of the multiplier at the $l$-th, $\forall l\in\mathbb{N}$, step is improved by $\mu^{(l)}=\frac{\mu_{-}^{(l)}+\mu_{+}^{(l)}}{2}$. As shown in Algorithm 1, at each step, the value of either $\mu_{-}^{(l)}$ or $\mu_{+}^{(l)}$ and the corresponding decision policy $\pi_{\gamma,\mu_{-}^{(l)}}$ or $\pi_{\gamma,\mu_{+}^{(l)}}$, respectively, are updated according to the cost constraint in (IV-B3) until a stopping criterion $\rvert\mu_{+}^{(l)}-\mu_{-}^{(l)}\rvert<\varepsilon_{\mu}$ is reached with the accuracy $\varepsilon_{\mu}$. Considering (IV-C) and (14), $h_{\gamma}(\mu)$ is a non-increasing function of $\mu$ in the unsettled region where the cost constraint has not yet been met. In this regard, the bisection method searches for the smallest Lagrange multiplier that guarantees the constraint. Also, one can show that $h_{\gamma}(\mu)$ denotes a Lipschitz continuous function with the Lipschitz constant as below
$\bigg{\lvert}\,C_{\gamma,\rm max}-\underset{N\rightarrow\infty}{\lim\sup}~{}\frac{1}{N}\,\mathbb{E}\bigg{[}\sum_{n=1}^{N}\lambda^{n}c_{\gamma}(\gamma_{n})\bigg{]}\bigg{\rvert}.$
Therefore, the bisection search converges to the optimal value of $\mu$ within finite steps [25, pp. 294].
@@
$\underset{N\rightarrow\infty}{\lim\sup}~{}\frac{1}{N}\,\mathbb{E}\bigg{[}\sum_{n=1}^{N}\lambda^{n}c_{\gamma}(\gamma_{n})\bigg{]}=C_{\gamma,\rm max}.$ (23)
Otherwise, the derived policy becomes randomized stationary in the shape of mixing two deterministic policies $\pi_{\gamma,\mu_{-}^{(l)}}$ and $\pi_{\gamma,\mu_{+}^{(l)}}$ with probability $\eta\in[0,1]$ [26], where
$\pi_{\gamma,\mu_{-}^{(l)}}=\underset{\mu\rightarrow\mu_{-}^{(l)}}{\lim}\pi_{\gamma,\mu},~{}~{}\pi_{\gamma,\mu_{+}^{(l)}}=\underset{\mu\rightarrow\mu_{+}^{(l)}}{\lim}\pi_{\gamma,\mu}$
in the $l$-th step, after which the bisection search terminates. Hence, we can write
$\pi_{\gamma}^{*}\leftarrow\eta\pi_{\gamma,\mu_{-}^{(l)}}+(1-\eta)\pi_{\gamma,\mu_{+}^{(l)}},$ (24)
which implies that the decision policy is randomly chosen as $\pi^{*}_{\gamma}=\pi_{\gamma,\mu_{-}^{(l)}}$ and $\pi^{*}_{\gamma}=\pi_{\gamma,\mu_{+}^{(l)}}$ with probabilities $\eta$ and $1-\eta$, respectively. In (24), $\eta$ is computed such that the condition in (23) is maintained.
@@ IV-D3 Complexity analysis
The value iteration approach in the inner loop is polynomial with $\mathcal{O}(\rvert\mathcal{A}_{\gamma}\rvert\rvert\mathcal{S}_{\gamma}\rvert^{2})$ arithmetic operations at each iteration. Besides, the bisection search takes $\lceil\log_{2}(\frac{\mu_{+}^{(0)}}{\epsilon_{\mu}})\rceil$ steps to reach the optimal Lagrange multiplier within the tolerance of $\varepsilon_{\mu}$, given the derived policy. Here, $\mu_{+}^{(0)}$ is the upper bound of the initial interval for the multiplier, with the lower bound fixed at zero. Thereby, the overall complexity of Algorithm 1, in terms of the number of arithmetic operations across both loops, can be calculated as follows
$\mathcal{O}\!\left(\frac{2\rvert\mathcal{S}_{\gamma}\rvert^{2}}{1-\lambda}\log\!\left(\frac{1}{1-\lambda}\right)\!\log\!\left(\frac{\mu_{+}^{(0)}}{\varepsilon_{\mu}}\right)\!\right)$
based on the fixed $\lambda<1$ (see [27] and [28]). The algorithm’s complexity increases as the state space expands, the initial interval for the multiplier widens, $\lambda\rightarrow 1$, and $\varepsilon_{\mu}\rightarrow 0$.
@@ V Monte Carlo Probability Distribution Estimation
Within the E-horizon, the SA does not make any decisions. Instead, it focuses on communicating updates at the highest possible rate while adhering to cost constraints. Once receiving these updates, the AA measures their usefulness, stores them in memory, and sends E-ACK signals for effective updates. The SA logs whether the E-ACK has been successfully received or not at every slot of the E-horizon. Finally, employing the received E-ACK signals at the SA and the measured updates’ usefulness at the AA, both agents perform their estimations.
@@ V-A Usefulness Probability of Received Updates
Picking the $j$-th, $\forall j\in\mathcal{J}$, outcome from the set $\hat{\mathcal{V}}$ (defined in Section III-A) that corresponds to the received update’s usefulness from the endpoint’s perspective at the $m$-th slot of the E-horizon, i.e., $\hat{v}_{m}$, the relevant estimated probability of that outcome is given by
@@ V-B Probability of the Mapped Target Usefulness
We assume that the mapped target usefulness, i.e., $\hat{v}_{\rm tgt}$, is a member of the set $\hat{\mathcal{V}}_{\rm tgt}=\{\vartheta_{j}\,\lvert\,j\in\mathcal{J}_{\rm tgt}\}$ with i.i.d. elements, where $\mathcal{J}_{\rm tgt}=\{1,2,\dots,\lvert\hat{\mathcal{V}}_{\rm tgt}\rvert\}$, and the probability of the $j$-th element is equal to $p_{\hat{v}_{\rm tgt}}(\hat{v}_{\rm tgt}=\vartheta_{j})$. Herein, as mentioned earlier, $p_{\hat{v}_{\rm tgt}}(\cdot)$ is the estimated pmf of the mapped target usefulness and obtained by
$p_{\hat{v}_{\rm tgt}}(\vartheta_{j})$ $=\sum_{e\in\{0,1\}}p_{\hat{v}_{\rm tgt}}\big{(}\vartheta_{j}\,\lvert\,\hat{E}=e\big{)}\mathrm{Pr}\big{(}\hat{E}=e\big{)}$ (26)
@@
where $\hat{E}_{m}$ indicates the E-ACK arrival status at the $m$-th slot of the E-horizon. Furthermore, to derive the conditional probability in (26), we first consider the successful arrivals of E-ACK signals such that
$p_{\hat{v}_{\rm tgt}}\big{(}\vartheta_{j}\,\lvert\,\hat{E}=1\big{)}$ $=\frac{\sum_{i\in\mathcal{I}}p_{\nu|\hat{E}=1}\big{(}\nu_{i}\,\lvert\,\nu_{i}\geq\vartheta_{j}\big{)}}{\sum_{j\in\mathcal{J}_{\rm tgt}}\sum_{i\in\mathcal{I}}p_{\nu|\hat{E}=1}\big{(}\nu_{i}\,\lvert\,\nu_{i}\geq\vartheta_{j}\big{)}}.$ (28)
@@
$p_{\hat{v}_{\rm tgt}}\big{(}\vartheta_{j}\,\lvert\,\hat{E}=0\big{)}$ $=\frac{\sum_{i\in\mathcal{I}}p_{\nu|\hat{E}=0}\big{(}\nu_{i}\,\lvert\,\nu_{i}<\vartheta_{j}\big{)}}{\sum_{j\in\mathcal{J}_{\rm tgt}}\sum_{i\in\mathcal{I}}p_{\nu|\hat{E}=0}\big{(}\nu_{i}\,\lvert\,\nu_{i}<\vartheta_{j}\big{)}}.$ (29)
@@ VI-A Setup and Assumptions
TABLE I: Parameters for simulation results
@@
Controlled transmission rate (effect-agnostic) $0.8$
Controlled query rate (effect-agnostic) $0.8$
@@ VI-B Results and Discussion
Fig. 4 illustrates the evolution of the average cumulative effectiveness over time for the push-and-pull model and different agent decision policies. As it is shown, applying the effect-aware policy in both agents offers the highest effectiveness, where the gap between its and the other policies’ performance increases gradually as time passes. Nevertheless, using the other effect-agnostic policies at either or both agent(s) diminishes the effectiveness performance of the system by at least $12\%$ or $36\%$, respectively. Particularly, if the SA and the AA apply the effect-aware and Markovian effect-agnostic policies, sequentially, the offered effectiveness is even lower than the scenario in which both agents use periodic effect-agnostic policies. One of the reasons is that the Markovian query raising misleads the SA in estimating the mapped target usefulness of updates. It is also worth mentioning that throughout the E-horizon from slot $1$ to $1\times 10^{5}$, the scenarios where the AA initiates effect-aware queries offer the same performance and better than those of the other scenarios. However, in the D-horizon, a performance gap appears and evolves depending on the applied policy at the SA. Therefore, by making effect-aware decisions at both agents based on estimations in the E-horizon and CMDP-based policies in the D-horizon, performance can be significantly improved.
@@
Figure 4: Evolution of the average effectiveness accumulated over time in the push-and-pull model.
@@
The bar chart in Fig. 5 shows the average rate of update transmissions from the SA and the consequent actions performed at the AA based on the initiated queries that result in the system’s effectiveness, as depicted in Fig. 4. Interestingly, the scenario where both agents apply the effect-aware policy has lower transmission and action rates compared to the other scenarios except the ones with Markovian effect-agnostic query policies. However, the number of actions that can be taken for those scenarios with Markovian queries is limited since most updates are received outside the action windows. Figs. 4 and 5 show that using effect-aware policies at both the SA and the AA not only brings the highest effectiveness but also needs lower update transmissions by an average of $11\%$, saving resources compared to the scenarios that have comparable performances. We also note that although effect-aware and periodic query decisions with effect-aware update transmission have almost the same action rate, the effect-aware case leads to more desirable effects or appropriate actions at the endpoint. This results in around $16\%$ higher average cumulative effectiveness, referring to Fig. 4.
@@
Figure 5: Average update transmission and query rates for different decision policies in the push-and-pull model.
( (a))
( (b))
( (c))
@@
Figure 6: Temporal evolution of the average cumulative GoE for the (a) push-and-pull, (b) push-based, and (c) pull-based models.
@@
Figs. 6 (a), 6 (b), and 6 (c) present the average cumulative GoE provided in the system over $5\times 10^{5}$ time slots for the push-and-pull, push- and pull-based communication models, respectively. The corresponding effectiveness at the endpoint for the primary model can be found in Fig. 4. The plots demonstrate that when both agents decide based on effect-aware policies, regardless of the update communication model, the highest offered GoE of the system is reached. However, in other scenarios, the performance of some policies exceeds those of others, depending on the update model. For instance, having effect-aware update transmission and Markovian queries shows a $2.52$ times higher average GoE for the push-based model than the push-and-pull one. This is because the push-based model has a larger action window.
@@
Besides, comparing Figs. 6 (a) and 6 (b), the reason that the provided GoE by the effect-aware decisions at both agents is $28\%$ lower for the push-based model compared to the push-and-pull is that the AA has to be available longer, which causes a higher cost. Also, with the pull-based model as in Fig. 6 (c), the average GoE within a period is less than the average cost since the AA is only available to act at query instants, significantly reducing the average GoE despite the high update transmission rate. In the pull-based model, however, applying the CMDP-based update transmission decisions at the SA can address this issue with a $16\%$ higher average GoE.
@@
The trade-off between the average effectiveness and the width of the action window, i.e., $\Theta_{\rm max}$, is shown in Fig. 7 for different decision policies. We see that the system cannot offer notable effectiveness with $\Theta_{\rm max}=1$, i.e., under the pull-based model. However, by expanding the action window width from $\Theta_{\rm max}=1$ to $10\,[\text{slot}]$, the average effectiveness boosts from its lowest to the highest possible value. Since all policies already reach their best performance before $\Theta_{\rm max}=10$, which indicates the push-based model, we can conclude that the push-and-pull model with a flexible action window is more advantageous than the push-based one with a very large fixed window. In addition, the scenario where both agents use effect-aware policies outperforms the others for $\Theta_{\rm max}>1$. It is worth mentioning that the performance of the scenarios where the AA initiates Markovian queries converges to the highest level for very large widths. The average cumulative effectiveness of these scenarios, as in Fig. 4, visibly rises for $\Theta_{\rm max}\geq 9$.
@@
Figure 7: Average effectiveness versus the action window width in the push-and-pull communication model.
@@
Figure 8: Average effectiveness as a function of the target effectiveness grade in the push-and-pull model.
@@
Fig. 9 depicts the average effectiveness obtained in the system through $5\times 10^{5}$ time slots versus the controlled update transmission rate for the push-and-pull model and various agent decision policies. Concerning Section II-B, this controlled rate is related to the effect-agnostic update transmission policies at the SA and denotes the expected number of updates to be communicated within the specified period. Therefore, the performance of the other policies should remain fixed for different controlled transmission rates. Fig. 9 reveals that increasing the controlled update rate increases the offered effectiveness when the SA applies the effect-agnostic policies. However, even at the highest possible rate, subject to the maximum discounted cost, using effect-aware policies at both agents is necessary to ensure the highest average effectiveness, regardless of the controlled transmission rate. As an illustrative example, when the AA initiates effect-aware queries, the effectiveness drops by an average of $38\%$ ($43\%$) if the SA transmits updates based on the periodic (Markovian) effect-agnostic policy instead of using the effect-aware one.
@@
Figure 9: Comparison between different policies with variable update transmission rates but fixed query rates for effect-agnostic policies.
In Fig. 10, the plot shows the same trend as Fig. 9, but this time it focuses on the changes in the average effectiveness versus the controlled query rate. Herein, the controlled query rate is dedicated to the scenarios where the AA operates under effect-agnostic policies. The results indicate that the average effectiveness rises via the increase of the controlled query rate for the periodic and Markovian policies. However, the increase in effectiveness is not significant for the latter one. Despite this, even with the highest rates, the effectiveness offered in the course of effect-aware queries is still higher than those of raising effect-agnostic queries. Therefore, the highest average effectiveness is achieved when both agents make effect-aware decisions, as depicted in Figs. 9 and 10.
Figure 10: Comparison between different policies with variable query rates but fixed update transmission rates for effect-agnostic policies.
@@
In the context of the decision-making problem $\mathcal{P}_{2}$ in (IV-B3), altering the maximum discounted cost, i.e., $C_{\gamma,\rm max}$, can impact the decisions made by each agent. To study this, we have plotted Fig. 11 for the push-and-pull model under different decision policies. The figure shows that the stricter the cost constraint, the lower the average effectiveness, irrespective of the decision policy. Also, for all cost constraints, the scenario in which both agents use effect-aware policy yields the best performance, whereas the other policies outperform each other under different cost constraints. Due to CMDP-based decisions, the gap between the effectiveness performance of the best scenario and those of the others increases as the constraint decreases until $C_{\gamma,\rm max}=0.1$, where the SA can transmit all updates, and the AA can initiate queries without restrictions.
@@
Figure 11: Average effectiveness achieved through different policies versus the maximum discounted cost in the push-and-pull model.
According to the CMDP modeled at either agent, model-based effect-aware policies heavily rely on the conditions of the update and/or E-ACK channels. To evaluate this impact, Fig. 12 illustrates the average effectiveness as a function of the channel erasure probability. For clarity, we assume $p_{\epsilon}=p_{\epsilon}^{\prime}$, where $p_{\epsilon}$ and $p_{\epsilon}^{\prime}$ denote the erasure probabilities in the forward communication and the acknowledgment links, respectively. At lower erasure probabilities, particularly when $p_{\epsilon}\leq 0.52$, employing effect-aware policies at both agents yields the highest average effectiveness. However, as channel conditions worsen, the performance of effect-aware decision-making gradually declines. Notably, when $p_{\epsilon}\geq 0.66$, the effectiveness of pulling updates using effect-aware policies drops to its lowest levels.
Figure 12: Interplay between average effectiveness and erasure probability for both the update and E-ACK channels in the push-and-pull model.